OpenStudy (anonymous):

Simplify 1/5(10x+15)-3(3x+4)

OpenStudy (anonymous):

is it:\[(1/5(10x+15)) - 3(3x+4)?\]

OpenStudy (anonymous):

11x+15

OpenStudy (anonymous):

what is the equation though?

OpenStudy (anonymous):

(1/5(10x+15))-3(3x+4)=(10x/5)+(15/5)-9x+12=(10x/5)+3+(45x/5)+12= (55x/5)+15= 11x+15

OpenStudy (anonymous):

alright then :)

OpenStudy (anonymous):

are you sure? because I got :\[= [1 - 15(3o0x^2 + 85x +60)]/5(10x+15) = [-450x^2 -1275x - 899]/5(10x+15)\] same denominator then subtract.

OpenStudy (anonymous):

15(30x^2) * sorry ^^"

OpenStudy (anonymous):

instead of 15(300x^2)

OpenStudy (anonymous):

i did mess up it should be a -45x/5 instead of the +45x/5 and a -12 instead of the +12 the answer should come out to -35x/5-8 or -7x-8

OpenStudy (anonymous):

are you sure? because that's not what I got, you multiply the other side to get the same denominator then subtract.